OpenStudy (anonymous):

Find the vertex and focus of y^2+6x=0

OpenStudy (sidsiddhartha):

6x+y^2=0 y^2=-6x This is an equation of a parabola that open leftwards. Its standard form: (y-k)^2=4p(x-h), (h,k)=(x,y) coordinates of the vertex For given equation: y^2=-6x vertex:(0,0) here 4*p=6 p=3/2 focus=(-3/2,0)